OpenStudy (anonymous):

lim of x of infinity 3xsquared -x+4 all over 2xsquared + 5x - 8 I have to show each step...help!

OpenStudy (anonymous):

There's really no good way to simplify that except to know that when there is a \[\lim_{x \rightarrow \infty}\]then you have to check the powers of the fraction. In this case, the highest power on top is 3 x squared, and the highest power on bottom is 2 x squared. Since you can cross off the lower powers, you're left with 3x squared over 2x squared which is 3/2.

OpenStudy (anonymous):

3/2

OpenStudy (anonymous):

what pjw said but you do it with your eyeballs

OpenStudy (anonymous):

First : divid all terms of the numerator & denomenator by the highest poer , i.e x^2

OpenStudy (anonymous):

however... we can write some steps if you like

OpenStudy (anonymous):

\[\frac{3x^2-x+4}{2x^2+5x-8}=\frac{\frac{3x^2}{x^2}-\frac{x}{x^2}+\frac{4}{x^2}}{\frac{2x^2}{x^2}+\frac{5x}{x^2}-\frac{8}{x^2}}\]

OpenStudy (anonymous):

you get \[\frac{3-\frac{1}{x}+\frac{4}{x^2}}{2+\frac{5}{x}-\frac{8}{x^2}}\]

OpenStudy (anonymous):

===> lim [( 3+4/x^2) /2+5/x^2-5/x^2] = (3+0)/(2+0+0)=3/2

OpenStudy (anonymous):

take the limit as x goes to infinity and see that everything goes to zero except the 3 and the 2 so limit is 3/2 but as you can see this is a colossal waste of time. you do it with your eyes. degree of numerator is the same as the degree of the denominator, take ratio of leading coefficient.

OpenStudy (anonymous):

nice! thanx!!!